### v4

# Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles with Reliable Explanations

Abstract
[…]
Pub-Guard-LLM: Detecting Fraudulent Biomedical Articles
with Reliable Explanations
Lihu Chen1, Shuojie Fu1, Gabriel Freedman1, Cemre Zor2, Guy Martin3, James Kinross1, Uddhav Vaghela1, Ovidiu Serban1, Francesca Toni1 1 Imperial College London, UK 2 Amazon Web Services, UK 3 National Health Service, UK
Figure 1: Workflow of the proposed Pub-Guard-LLM (Details in Section 4)
1 Introduction
A considerable proportion of published scientific articles has been found to involve fraudulent practices. It is estimated that between 1.5 and 2% of all scientific papers published in 2022 closely resemble content produced by paper-mills Noorden (2023), and about 8,000 papers were retracted from Hindawi journals due to their fraudulent origins in paper mills in 2023 Parker et al. (2024). Even more concerning, generative AI tools including those built on large language models (LLMs) can produce highly convincing fraudulent articles that have the potential to bypass existing detection mechanisms Májovskỳ et al. (2023); Kendall and Teixeira da Silva (2024). This indicates that fraud in this area is constantly adapting and learning to circumvent existing detectors Májovskỳ et al. (2023); Perkins (2023).
Fraudulence of biomedical publications is more pronounced than in other fields Grieneisen and Zhang (2012); Bik et al. (2016), with a recent finding revealing that over a fifth of newly published medical articles are at least partially fraudulent Sabel et al. (2023). Moreover, the problem of fraudulent article detection is not confined to a specific set of offending author institutions. In fact, recent high-profile cases have uncovered academic fraud from credible sources such as Harvard Medical School Johnson (2024). This widespread trend poses a serious threat to public health, as misleading medical research can directly influence clinical decisions, leading to ineffective or even harmful treatments for patients.
Prior attempts addressing this challenge mainly rely on relatively primitive heuristic techniques Parker et al. (2022); Shepperd and Yousefi (2023); Feng et al. (2024), including some based on argument quality Freedman and Toni (2024). While the threat of fraudulent articles is likely to grow in the future, the NLP community has yet to give this issue the attention it deserves. This lack of focus is partly due to two key challenges. First, there is no standard benchmark dataset for evaluating and comparing various fraud detection systems. Second, there are no open-source systems leveraging on LLMs specially designed for this task.
To address these challenges, we introduce PubMed Retraction, the first large-scale open-source benchmark for fraud detection in biomedical research, including over 11K real-world articles. Our benchmark is designed to consider the diversity across various types of academic misconduct, rather than focusing solely on specific issues like plagiarism or machine-generated text. The main goal of this benchmark is to flag potentially fraudulent articles with minimum information [^1] Complex features, such as tables and images, are not the focus of the current version of our benchmark.. Also, we release the first LLM-based system, Pub-Guard-LLM, dedicated to the task of fraud detection. Pub-Guard-LLM can be deployed in three application modes (see Figure 1): Vanilla Reasoning, Retrieval-Augmented generation, and Multi-Agent Debate, each making use of external knowledge and fine-tuning and each returning (textual) explanations of predictions. We show that Pub-Guard-LLM can significantly outperform baselines while generating more reliable explanations, by being more relevant to the explanandum Kotonya and Toni (2024) and more coherent Kotonya and Toni (2020). The three application modes provide options for users to focus on depending on their needs, balancing key factors such as precision and recall of predictions, reliability of explanations, and inference speed. This adaptability makes Pub-Guard-LLM highly versatile, accommodating a wide range of use cases in fraudulent article detection.
With the introduction of the first publicly available benchmark and an open-source LLM specifically designed for fraudulent article detection, our work aims to serve as a foundational step for future research in this field. However, while technological solutions can help address this issue, the root of the problem lies within academic evaluation systems themselves. Researchers face immense pressure to publish to secure funding and career advancements. As long as these systems prioritize metrics and publication counts over genuine scientific contributions, the arms race between fraud makers and detectors will persist.
2 Related Work
The increasing scientific fraud presents a serious challenge to both the research community and society at large. Previous studies addressing this challenge primarily focus on rule-based heuristics, based on analyzing indicators such as author affiliations Sabel et al. (2023), citation patterns Shepperd and Yousefi (2023); Feng et al. (2024), journal impact factors and author networks Pérez-Neri et al. (2022). Recently, traditional machine learning techniques Dadkhah et al. (2023); Cabanac et al. (2022) as well as BERT-based models Freedman and Toni (2024) have been proposed to help with fraudulent article detection.
However, the rapid growth in the number of fraudulent articles has not been matched by dedicated attention within the NLP field  Byrne et al. (2024). Currently, there is an urgent need for more research into two key aspects. (1) Benchmarks. Existing studies generally employ Retraction Watch The Center for Scientific Integrity (2018), a blog that monitors and reports on retractions of scientific papers, to analyze retraction behaviors Shepperd and Yousefi (2023); Byrne et al. (2022); Freedman and Toni (2024). The Retraction Watch provides a collection of retracted records, but it does not constitute a benchmark on its own for comparing existing methods. While benchmarks exist for specific fraud types, such as plagiarism Wahle et al. (2021) and machine-generated text Mosca et al. (2023); Abdalla et al. (2023); Liyanage et al. (2022), there is currently no standardized benchmark including diverse fraud types using real-world articles. (2) Open-Source Tools. While some systems for detecting fraudulent articles exist, they are often proprietary and controlled by commercial entities that are reluctant to share their technologies Christopher (2021). Even though some LLM-based applications have been developed to address news misinformation Cao et al. (2024) and plagiarism Wahle et al. (2022), there is a need for open-source LLMs specifically tailored to scientific fraud detection. Such models would enhance transparency and foster collaboration in the fight against scientific fraud.
Note that the development of open-source tools alone is not sufficient; effective fraud detection systems must also be transparent and interpretable. The necessity for systems to provide explanations for their outputs in tasks such as fact-checking Kotonya and Toni (2020) and misinformation detection Schmitt et al. (2024) is well-acknowledged. While previous research has explored incorporating explainable AI to explain decision-making processes Dalal et al. (2024); Kotonya and Toni (2024), the study of explainability in fraud detection systems remains limited.
General Cancer
Dataset Train Validation Test Breast Lung Ovarian Colorectal
Sample 10,000 500 500 300 300 292 300
Fraudulence Rate 24.5% 25.6% 22.4% 25.0% 38.7% 38.0% 33.0%
High Profile Rate 34.0% 35.9% 33.9% 38.7% 22.4% 29.7% 28.3%
Table 1: Basic statistics of Pubmed Retraction. High Profile Rate refers to the articles deemed fraudulent despite their meta-data being ranked high (details in Appendix A).
3 PubMed Retraction Benchmark
To address the lack of a standardized benchmark for evaluating and comparing fraudulent article detection methods, we curate a large open-source benchmark for biomedical articles, PubMed Retraction, comprising over 11K real-world biomedical articles, including both legitimate and retracted publications. We select the biomedical domain due to the higher prevalence of fraudulent publications therein than in other fields  Grieneisen and Zhang (2012); Bik et al. (2016), as well as the significant impact of this type of fraud in patient safety, clinical decision-making, and public trust in medical research.
To construct PubMed Retraction, we use the PubMed database Wheeler et al. (2007) and retrieve articles using the Unix command-line tool EDirect Kans (2024). To ensure relevance and avoid outdated data, we focus on articles published between 2000 and 2025. Our goal is to promote the use of minimal textual information for fraud detection, so we design the benchmark to include only title, abstract, and metadata while excluding more complex features such as full text and images. Specifically, each article in PubMed Retraction is structured to include seven key attributes: Pubmed ID, Title, Abstract, Authors, Affiliations, Journal, Is Fraudulent, with the latter a binary indicator (Yes or No). An article is labeled as fraudulent if the keyword “retracted” appears in its publication type attribute. Note that while retracted articles are not always explicitly fraudulent, they often contain errors, misconduct or unreliable findings Grieneisen and Zhang (2012), indicating potential fraud that needs to be surfaced by fraudulent article detection. To maintain a realistic distribution, we include a lower number of fraudulent articles than legitimate ones, which is aligned with real-world cases. Overall, the benchmark consists of 11,192 articles.
To evaluate generalization capabilities of models, we partition articles based on keywords to create two subsets as described below. This partitioning strategy is designed for a fair assessment of a model’s ability to handle out-of-distribution samples, a critical aspect for real-world applications where models encounter unseen data. (1) General Set. This set comprises articles covering various diseases and is used for model fine-tuning. It is further split into Train, Validation, and Test subsets. (2) Cancer Set. This set consists exclusively of articles related to four types of cancer: Breast, Lung, Ovarian, and Colorectal. Articles of these four categories are excluded from the General Set, which means that these cancers are unseen diseases during fine-tuning and validation. Additionally, our benchmark includes high-profile fraudulent cases, which makes this task particularly challenging (details in Appendix A).
Formally, ignoring the partitions, PubMed Retraction can be represented by $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}$ is an article (title, abstract and metadata) and $y_{i}$ is the binary label. The basic statistics of Pubmed Retraction are given in Table 1.
4 Pub-Guard-LLM
Pub-Guard-LLM is an end-to-end system for fraudulent article detection. Figure 1 provides an overview: First, the input data is enriched with external knowledge. Second, explanations to support article labels (fraudulent or not) are obtained by distilling a strong LLM (Teacher Model) via zero-shot prompting. Finally, using the article attributes and external knowledge as inputs, and distilled explanations with article labels as outputs, we fine-tune an LLM to create Pub-Guard-LLM. We propose three application modes for Pub-Guard-LLM, resulting in different performances as concerns F1, recall, reliability of explanations, and inference speed.
4.1 External Knowledge Workflow
To complement the seven input attributes (see Section 3) and incorporate task-specific knowledge, we augment external knowledge into Pub-Guard-LLM, improving LLM reasoning—a widely adopted strategy in NLP Pan et al. (2024). For this, we identify three fundamental factors to assess an article’s legitimacy: (1) author credibility, (2) affiliation reputation, and (3) journal reputation, and use the relevant external knowledge to enrich the input.
Author Credibility.
In order to evaluate the credibility of authors, we use the Semantic Scholar database Lo et al. (2019), which contains millions of articles with extensive metadata and citation information, as an external knowledge source. To quantify author reputation, we use the h-index. Since LLMs are not inherently sensitive to numerical values, we map h-index scores into five reputation levels (very low, low, medium, high, very high ) using a predefined piecewise function (see Appendix A), e.g., Geoffrey Hinton (author h-index: 188, very high).
Affiliation Reputation.
As an indicator of the reputation of an author’s affiliation, we use the average citation count per institution, determined with OpenAlex Priem et al. (2022) as external knowledge source. This is a fully open index of scholarly works, which provides comprehensive data on the total publications and citations of institutions. The calculated average citation count is then mapped into five reputation levels (very low, low, medium, high, very high), e.g., Harvard University (institution average citation: 64, very high).
Journal Reputation.
We use the ImpactFactor library222%https://impact-factor.readthedocs.io/en/latest/ to obtain the Journal Citation Reports (JCR) partition, which serves as an indicator of a journal’s reputation. To enhance interpretability, we assign a human-readable categorized label (low, medium, high or very high) to each journal based on its JCR ranking, e.g., Nature (journal JCR: Q1, very high).
The knowledge retrieved from external resources is integrated into the benchmark $\mathcal{D}$ for reducing duplicated queries, formally resulting in $\mathcal{D}^{\prime}=\{(x_{i},k_{i},y_{i})\}_{i=1}^{N}$ where $k_{i}$ is the external information for $(x_{i},y_{i})\in\mathcal{D}$. Note that if the inquired knowledge is missing from the external knowledge sources, we use, as $k_{i}$, the keyword “null” to indicate the absence. We choose to do so rather than omitting the field entirely, as we found it does generally indicate a lower article quality and improves the downstream performance.
4.2 Explanation Distillation
The PubMed Retraction benchmark provides only a binary label (fraudulent or not) as a target. To construct a model that classifies articles accurately while offering reliable explanations, it is essential to incorporate either human-annotated or machine-generated explanations into the ground truth. Given the high cost of human annotation, we opt for the latter as a more scalable and efficient alternative.
Recent research shows that distilling reasoning capabilities from larger (language) models can significantly enhance the performance of smaller models with fewer learning examples Li et al. (2022); Hsieh et al. (2023); Shridhar et al. (2023): teacher model rationales provide richer insights into why an input is mapped to a specific output label, capturing task-relevant knowledge missing from the input. We use a teacher model to generate explanations for each article via prompting, resulting in $\mathcal{D}^{\prime\prime}=\{(x_{i},k_{i},e_{i},y_{i})\}_{i=1}^{N}$ where $e_{i}$ is the corresponding distilled explanations with regard to $(x_{i},k_{i},y_{i})\in\mathcal{D}^{\prime}$.
Note that instruction-aligned LLMs tend to classify all articles as not fraudulent, likely to avoid controversial responses. This behavior arises because many LLMs, when fine-tuned for alignment and safety, may become overly conservative and ignore contextually relevant information Röttger et al. (2024). To mitigate this issue, it is crucial to assign a clear stance to the teacher model. For example, when evaluating a retracted article, we include the prompt: “You are tasked with providing explanations and making a firm case for why it has issues and should be retracted. You should have a clear stance.”. This ensures strong, rationale-driven explanations instead of the default neutral or overly cautious responses (see Table A2 in Appendix for an example prompt).
4.3 Fine-Tuning
We leverage the samples in PubMed Retraction augmented with the external information and distilled explanations ($\mathcal{D}^{\prime\prime}$) to fine-tune Pub-Guard-LLM. To encourage Pub-Guard-LLM to learn both labels and explanations, we introduce a multi-task learning objective. Suppose we have a base model $f_{\theta}$ able to generate label $y$ and explanation $e$ given article $x$ with external information $k$ , i.e., $f_{\theta}(x,k)=(y,e)$. Then, our goal is to optimize $\theta$ for this task.
The model is fine-tuned on $y$ to understand how to identify fraudulent articles via a binary classification task with:
$\mathcal{L}_{\text{cls}}=\frac{1}{N}\sum_{i=1}^{N}\ell\big{(}f_{\theta}(x_{i},k_{i}),y_{i}\big{)}$ (1)
where $\ell$ is the cross entropy between the ground truth and the predicted label. To encourage the model to output explanations behind the decision, we introduce a next-token prediction task with:
$\mathcal{L}_{\text{explanation}}=\frac{1}{N}\sum_{i=1}^{N}\ell\big{(}f_{\theta}(x_{i},k_{i}),e_{i}\big{)}$ (2)
where $\ell$ is the averaged cross entropy between the generated tokens and the distilled explanation from the teacher model.
During fine-tuning, the model learns to predict both label $y$ and explanation $e$ given an article:
$\mathcal{L}=\mathcal{L}_{\text{cls}}+\lambda\mathcal{L}_{\text{explanation}}$ (3)
where $\lambda$ is a hyper-parameter that controls the balance between the two objectives. This learning process helps Pub-Guard-LLM acquire knowledge beyond simple binary labels, which results in strong performance with well-reasoned explanations.
Figure 2: Debate application mode for Pub-Guard-LLM
4.4 Application Modes
After fine-tuning, the model $f_{\theta}$ can be leveraged to identify fraudulent articles while providing explanations. To showcase its usability, we introduce three distinct application modes (see Appendix B.3 for input-output examples for each mode).
Vanilla Reasoning.
This refers to reasoning without requiring additional in-context learning samples (Zero-Shot Reasoning), which makes predictions solely based on the article: $f_{\theta}(x,k)$.
Retrieval-Augmented Generation (RAG).
$f_{\theta}(x,k)$ does not store all articles within its parameters, and, in some cases, an input article’s main argument may lack support from existing publications. This suggests that the article’s findings could be controversial and potentially fraudulent. To address this, we retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$ by passing them as input into the model: $f_{\theta}(x,k,\mathcal{A})$ (see details in Appendix B.2).
Debate.
Recent research has shown that Multi-Agent Debate can enhance both factual accuracy and explainability in LLMs Liang et al. (2023); Du et al. (2023); Freedman et al. (2024). The core idea involves multiple agents presenting their arguments, while a judge oversees to reach a final decision.
To generate the debate process in our setting (Figure 2), we train two specialized Pub-Guard-LLMs: (1) a support reviewer $f_{\theta}^{s}$, trained on legitimate articles, generates supporting arguments $arg^{+}$; (2) an attack reviewer $f_{\theta}^{a}$, trained on retracted articles, generates attacking arguments $arg^{-}$. Thus, given input article $(x,k)$, the support reviewer $f_{\theta}^{s}$ consistently provides arguments in favor of the article while the attack reviewer $f_{\theta}^{a}$ presents counterarguments highlighting potential issues. Following this debate, we introduce a meta-reviewer $f_{\theta}^{meta}$ that is responsible for making the final decision based on the arguments presented by both reviewers. To train the meta-reviewer, we use a teacher model to generate debate-based explanations $\hat{e}$, which are then used to fine-tune $f_{\theta}^{meta}$ by reusing the multi-task learning in Equation 3. Overall, the debate-based detection framework is formulated as: $f_{\theta}^{meta}(x,k,arg^{+},arg^{-})$, where the meta-reviewer makes a decision by considering both the supporting and attacking arguments.
Model $|\theta|$ Test Breast Lung Ovarian Colorectal Avg
Fine-Tuning Binary Classifiers
SciBERT 110M 60.9 56.1 74.0 67.2 71.5 66.0
BioLinkBERT 340M 58.1 57.7 75.4 71.0 72.2 66.9
Gatortron 345M 57.6 58.6 71.2 70.9 67.0 65.1
Llama-3.1 8B 60.9 63.2 76.5 72.3 68.3 68.2
Bio-Llama 8B 56.4 61.7 63.9 66.7 61.6 62.1
BioLinkBERT + PEARL 210M 67.6 69.6 79.0 70.5 69.8 71.3
Gatortron + PEARL 445M 66.3 69.3 78.3 76.2 70.8 72.2
Few-Shot Reasoning
Llama-3.1-Instruct 8B 29.8 20.0 46.2 43.3 43.6 36.6
OpenScholar 8B 11.5 18.8 20.3 19.9 14.1 16.9
Bio-Llama 8B 37.7 40.7 56.7 59.3 47.9 48.5
PMC-Llama 13B 36.0 36.4 56.7 23.3 25.8 35.6
Ours
Pub-Guard-LLM (Vanilla) 8B 70.7 72.8 79.2 78.3 75.6 75.3
Pub-Guard-LLM (RAG) 8B 69.8 76.1 82.3 79.5 76.4 76.8
Pub-Guard-LLM (Debate) 8B 70.1 71.5 81.1 78.0 75.7 75.3
Table 2: Performance in detecting fraudulent articles. We report average F1 scores over three seeds.
Test Breast Lung Ovarian Colorectal Avg
Vanilla 84.8 88.5 89.2 85.9 69.7 83.6
RAG 66.9 82.7 87.9 85.6 81.8 81.0
Debate 90.7 92.2 94.6 89.9 75.9 88.7
Table 3: Recall (average over 3 seeds) across modes
5 Experimental Set-Up
5.1 Baselines
We compare our approach to the following:
(1) Fine-tuning Binary Classifiers. This method appends a classification head—a linear layer—on top of a language model to handle fraud detection. However, this method can only output binary labels (no explanations). The title, abstract, and metadata are concatenated using a special token, [SEP], and this combined textual input is fed into the classifier. Specifically, we use language models: SciBERT Beltagy et al. (2019), BioLinkBERT Yasunaga et al. (2022), Gatortron Yang et al. (2022), Llama-3.1 Dubey et al. (2024) and Bio-Llama ContactDoctor (2024). We also find that introducing additional models to represent metadata separately can be beneficial. Thus, we apply a lightweight short-text model, PEARL Chen et al. (2024), to represent metadata and obtain embeddings for authors, affiliations, and journal. We continue to use the original language model to represent the title and abstract. Subsequently, these embeddings are concatenated and fed into the classification head. We denote the resulting variant of language model X as “X + PEARL”.
(2) Few-shot Reasoning. LLMs are generally known as few-shot reasoners, meaning they can perform powerful reasoning with few learning examples Brown et al. (2020); Kojima et al. (2022). Compared to binary classifiers, LLMs can offer answers with explanations. We provide LLMs with a certain number of learning examples and prompt them to obtain binary answers with explanations. Here, we use the following instruction-aligned LLMs: Llama-3.1-Instruct Dubey et al. (2024), OpenScholar Asai et al. (2024), Bio-Llama ContactDoctor (2024) and PMC-Llama Wu et al. (2024).
5.2 Implementation Details
All approaches are implemented with PyTorch Paszke et al. (2019) and HuggingFace Wolf et al. (2020). We use Amazon EC2 g5.12xlarge instances with 4$\times$24 GiB for all experiments. The Teacher model used for distilling explanations is Mistral-Large MistralAI (2024) on Amazon Bedrock. Our model fine-tunes Llama-3.1-8B Dubey et al. (2024) using LoRA Hu et al. (2021) adapters (r=128, lora_alpha=128, lora_dropout=0.1). We train on the train set in the General partition for one epoch and AdamW 8-bit optimizer with a learning rate of 1e-4, batch size 4, gradient accumulation of 4 and the $\lambda$ is 1.
All baseline binary classifiers are fine-tuned with three different seeds: we report their performance based on the best validation set score. For few-shot reasoning, we set the sample size to zero (zero-shot reasoning) since adding contextual examples provides only minimal benefits (1–2 percentage points).
6 Results
6.1 Overall Performance
Table 2 gives the results of the comparison of Pub-Guard-LLM with the baselines.
First, we note that Pub-Guard-LLM, across all three modes, consistently outperforms the baselines. The performances across the three modes do not exhibit significant differences, but the RAG mode achieves the best score on average, which shows that introducing relevant articles is beneficial to verify the claim quality in abstracts. On the other hand, a key advantage of the debate mode is its notably higher recall, as shown in Table 3. Achieving high recall is crucial in the fraud detection task, as its primary goal is to flag suspected fraud and provide warnings to end-users.
Second, BERT-based models prove to be cost-effective while achieving competitive performance comparable to larger counterparts in domain-specific tasks, which aligns with prior findings Lehman et al. (2023); Chen and Varoquaux (2024). For instance, Gatortron + PEARL ranks second overall. Additionally, leveraging top-layer representations of decoder-only models to train classifiers appears suboptimal, as it does not perform better than BERT-based models. Existing LLMs struggle in the few-shot setting, suggesting that domain adaptation is necessary for this task. Note that we do not report the performance of GPT-4o and Mistral-Large because these large instruction-aligned models frequently classify all articles as legitimate to avoid controversial conclusions.
In summary, all three modes of Pub-Guard-LLM effectively detect fraudulent articles and each offers distinct advantages. The Vanilla Reasoning mode is efficient and user-friendly, the RAG mode achieves the best overall performance, and the Debate mode excels in recall. This variability makes Pub-Guard-LLM highly versatile, accommodating a wide range of possible user needs in fraudulent article detection. Additionally, we conduct an ablation study to validate the effect of each component in our system, such as base LLM, Teacher Model, External Knowledge. The results are in Table A1 in the Appendix.
Figure 3: Explanation reliability using LLM-as-a-Judge (average over three runs)
(a) Author Credibility
(b) Affiliation Reputation
(c) Journal Reputation
Figure 4: Performance on the combined test sets (test plus cancer) of Pub-Guard-LLM (Vanilla) depending on the metadata levels. The horizontal line in each figure represents the F1 score achieved using corresponding heuristic features (see details in Appendix C.1).
Figure 5: Comparison of explanation reliability and label-explanation agreement in the user study (see details in Appendix C.2).
6.2 Explanation Reliability
Compared to BERT-based models, a distinct benefit of Pub-Guard-LLM is the explanations. To evaluate the reliability of LLM-generated explanations, we consider two dimensions introduced, for different settings, in prior work Kotonya and Toni (2020, 2024); Valentino and Freitas (2024). (1) Relevance. A relevant explanation should highlight the particular features and patterns within the article that the model identified as indicative of fraud or not. The model should only reference information that is present in the article, avoiding hallucinations. (2) Coherence. This refers to the logical consistency and clarity of the explanation provided by the model for its predictions. Coherence ensures that the explanation flows logically, thereby enhancing the user’s understanding of why an article is classified as fraudulent or not. To measure these properties, we adopt the LLM-as-a-Judge paradigm, which has been shown to align with both controlled and crowdsourced human preferences when using strong LLM judges like GPT-4  Zheng et al. (2023). We design prompts (see Appendices A4 and A5) for these two dimensions and ask GPT-4o to assess the provided explanations, which assigns a score between 1 and 10. The results, presented in Figure 3, show that our explanations outperform those of other LLMs, with the exception of OpenScholar. While OpenScholar demonstrates strong explanation quality, its fraud detection capability is significantly lower than other LLMs, with an average F1 score of just 16.9 (see Table 2). Also, note that the debate mode can reduce irrelevant explanations by achieving higher relevance scores than the another two modes. These findings validate that Pub-Guard-LLM not only achieves powerful performance but also provides reliable explanations.
6.3 Impact of Metadata Levels
We explore whether the external knowledge we use to assess levels of the metadata can provide useful heuristics for fraud detection. To test this, we compare Pub-Guard-LLM (Vanilla) to heuristic methods, e.g., using the credibility of authors as indicators (see Appendix C.1). The results are shown in Figure 4. First, we observe that relying solely on heuristic features is insufficient for fraud detection since Pub-Guard-LLM significantly outperforms heuristics on average. Second, prestigious researchers, institutions and journals should not be blindly trusted, and these high-profile frauds are harder to detect, as shown by the performance declines on high and very high groups in Figure 4a and Figure 4b. Additionally, null is an indicator of low-level credibility and Pub-Guard-LLM is robust against missing information. In summary, high-profile cases pose a significant challenge to fraud detection. In the future, we plan to investigate features beyond metadata such as images.
6.4 User Study
We conducted a pilot user study to evaluate whether the outputs of Pub-Guard-LLM align with users’ needs. To achieve this, we designed questionnaires covering our three application modes and gathered feedback from three experienced doctors (see details in Appendix C.2). The questionnaire focused on two key aspects: (1) Reliability: is the explanation coherent and relevant? (2) Agreement: do users agree with the predicted label and the explanation? Figure 5 presents the user study results. The feedback suggests that doctors find our model’s explanations reliable, but there is a discrepancy between expert opinions and machine-generated predictions regarding the debate mode. However, due to the limited number of participants, it is difficult to determine which application mode performs better.
7 Conclusion
In this work, we contribute to fraudulent academic articles by introducing a publicly available benchmark and the first open-source LLM-based system tailored for this task. Our Pub-Guard-LLM not only surpasses existing baselines but also delivers reliable, well-grounded explanations. We aim to raise awareness of academic fraud within the research community and drive further advancements to tackle this growing issue.
We emphasize the need for broader efforts to combat academic fraud in the following key areas. First, high-profile publications should not be blindly trusted, as fraudulent publications may still exist within them. Detecting such cases requires more than analyzing metadata and abstracts, and future studies should also investigate full texts and images to identify frauds. Second, as we develop technological solutions to detect fraud, entities like paper mills continuously improve their methods to circumvent detection. This ongoing arms race requires continuous adaptation of detection techniques. Finally, addressing scientific fraud fundamentally depends on human and societal factors. Researchers should adhere to scientific integrity, and our academic evaluation system should prioritize genuine value creation over mere quantitative metrics.
Limitations
While our study makes contributions to fraudulent article detection, we acknowledge two primary limitations: (1) Catastrophic forgetting. Pub-Guard-LLM is a task-specific model that may not perform well in general-purpose reasoning due to catastrophic forgetting McCloskey and Cohen (1989)—a phenomenon where a neural model loses previously acquired knowledge upon learning new information. (2) Limited participants in user study. Our user study involved a small number of participants, which limits the ability to systematically compare the three application modes. A more comprehensive study with a larger participant pool is necessary for thorough evaluation.
Appendix A Details of Benchmark and External Knowledge
Author Credibility
To categorize authors based on their reputation, we define a piecewise function that maps h-index values into discrete reputation levels. This approach ensures that LLMs can interpret author credibility effectively by using textual labels instead of raw numerical values. The categorization is as follows:
[…]
Affiliation Reputation
To assess the reputation of an author’s affiliation, we use the average citation count per institution, which serves as a key indicator of an institution’s scholarly influence. Likewise, we map these values into five distinct reputation levels:
  • Developing Institution (very low): $0\leq$ average citations $\leq 5$
  • Emerging Institution (low): $6\leq$ average citations $\leq 15$
  • Established Institution (medium): $16\leq$ average citations $\leq 30$
  • Reputable Institution (high): $31\leq$ average citations $\leq 45$
  • World-Class Institution (very high): average citations $>45$
Journal Reputation
To assess the reputation of a journal, we use the JCR, which classifies journals into four quartiles based on their impact factor. To enhance interpretability, we map JCR quartiles into human-readable reputation levels:
  • Top-Level Journal (very high): JCR Q1
  • High-Level Journal (high): JCR Q2
  • Moderate-Level Journal (medium): JCR Q3
  • Low-Level Journal (low): JCR Q4
Definition of High Profile
[…]
Missing Rate
In our analysis of the PubMed Retraction dataset, we observed that certain external information may be absent from specific databases. For instance, a journal might not be indexed by the JCR. To quantify this, we computed the missing rates for each feature within our dataset:
  • Author Credibility: 10.6% missing
  • Affiliation Reputation: 33.4% missing
  • Journal Reputation: Journal Indexing: 41.5% missing
Appendix B Details of Pub-Guard-LLM
B.1 Prompts
Here, we list the prompts used in this work: (1) Explanation Distillation, Table A2; (2) Prompt of Detecting Fraudulent Articles, Table A3; (3) Prompt of Relevance Evaluation, Table A4; (4) Prompt of Relevance Coherence, Table A5.
B.2 RAG Mode
We retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$. In our experiment, the value $l$ is 5 and the PubMed database is MedRAG PubMed [^3] https://huggingface.co/NeuML/pubmedbert-base-embeddings with 3.5M articles. The retriever here is PubMedBERT and we use cosine similarity to compute the semantic relatedness.
B.3 Output Examples
We provide an input-output example for each application mode: Vanilla Reasoning (Table A6); RAG (Table A7); Debate (Table A8).
Appendix C Details of Experiments
C.1 Heuristic Features
Since we incorporate external knowledge to represent metadata levels, these fine-grained features can serve as indicators for fraud detection. To leverage this, we design three heuristics based on different metadata attributes. The intuition is that articles from prestigious researchers or well-known institutions are generally perceived as more trustworthy. To classify articles as fraudulent or not, we establish a threshold-based approach using metadata credibility scores. For example, if an article’s highest author credibility falls within [medium, high, very high], we classify it as legitimate. Author credibility and affiliation reputation are categorized into six levels (including null). We define the three lower levels as indicators of fraudulence and the three higher levels as indicators of legitimacy. Journal reputation, which consists of five levels, is categorized similarly: the three lower levels indicate potential fraud, while the higher levels suggest legitimacy.
C.2 User Study
In the user study, we prepared three questionnaires to evaluate the Vanilla Reasoning, RAG and Debate modes independently. Using the Vanilla Reasoning mode, we generated explanations for 300 papers. We randomly selected five fraudulent papers and five legitimate papers from the result and again randomized the order of the paper to create the questionnaire. For each paper, we added 1) the information of the paper as it was in the model prompt; and 2) the explanations generated by the model. Three questions were asked for each paper: 1) whether the explanation was coherent and relevant; 2) to infer the predicted label from the given explanation; and 3) whether the user agreed with the actual prediction. The same process was followed for the other two questionnaires for the Debate and RAG models, except that, for the Debate model, the information of a supporting viewer and an attacking reviewer was added to the paper description. In total, 30 distinct papers were selected for evaluation by three clinicians.
Given the disagreement between human experts and model-generated outputs regarding the debate mode, we present the first three comments below. We hope these critical insights will be valuable for future studies.
  • I still don’t understand what or who the "attacking reviewer" is. Is this a real world person who has completed a peer review? I feel that labelling this as fraudulent without more detailed evidence of actual fraud is not correct. This is very different from this simply being "not very good" science. I don’t think it is fair to effectively discredit an author simply because they have a low H factor. Everyone has to start somewhere. Moreover, the interpretation does not provide any evidence at all that the authors have either falsified data, that they have a track record of falsifying data, that the authors are affiliated with authors suspected activity or that they have links to governmental or state actors that may want to push falsified data. There is no comment on funding or conflicts of interest which I would expect to be in this analysis and which is missing. The critique argues that "specific" data is missing from the abstract, but then it does not state what this is and what would be required for the abstract to be robust. In summary, this LLM suggests this is research fraud, when as presented, the data suggest that more simply the quality of the abstract is low. This is an important distinction, that could lead to legal proceedings.
  • I think this LLM review is a false positive. I think the criticism of the abstract is not founded. They have performed in vivo and in vitro validation studies, and knock out experiments. While detail is missing, this is an overly harsh critique of a piece of basic science, that is not pretending to be clinical data. Again, what data is missing form the author’s institutions? It doesn’t actually state which part of this data is important and why. We know they are from a reputable Japanese university, which seems enough to me. Again, to call in to question the credibility of the scientist is libellous. This would need much stronger evidence of criminal or fraudulent behaviour which is clearly missing. So, I think this is just a case of a not very well written abstract, that has been mis labelled as fraud.
  • The LLM is written in an overly cautious manner, which makes me lack trust in it. Firstly, there is a paradox in its explanation. It states the authors have a high H index, but the institution is emerging… so which are we to judge as being more important? This to me is biasing smaller organisations that may produce outstanding work. We need this model to identify EXCELLENCE in science and to promote it where it exists irrespective of the institution. The factors these models are basing their analysis on e.g. H factor, institutional reputation and author details are not important. I want to understand patterns of publishing behaviours, previous track records of fraud, history of COIs… etc. etc.. and most importantly a robust analysis of the QUALITY of the science. The LLM really provides no insights on this.
C.3 Ablation Study
To validate the effect of different components in our methodology, we either vary or remove them and observe their impact on model performance, as measured by F1@Val in Table A1. First, we observe that fine-tuning is the most important component, as its removal causes a dramatic performance drop. This indicates it is necessary to adapt LLMs to this task of fraud detection. External knowledge is very beneficial, and the removal leads to a significant decrease. Meanwhile, other components can provide meaningful benefits, which confirms the effectiveness of our methodology.
F1@Val $\Delta$
Pub-Guard-LLM (Vanilla) 69.5 -
Varying the Base LLM
Bio-Llama 68.6 -0.9
Mistral 67.8 -1.7
Varying the Teacher Model
GPT-4o 68.3 -1.2
Llama-70B 66.4 -3.1
Removing Components
without Fine-tuning 29.9 -39.6
without Explanation 68.5 -1.0
without External Knowledge 59.8 -9.7
Table A1: Ablation study results
Prompt of Explanation Distillation
This following article is a {label} article.
You are tasked with providing explanations and making a firm case for why it has issues and should be {label}.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
Based on the provided information, give a very short and concise explanation why it should be {label}
Here are some factors to consider:
(1) the reputation of the journal (whether it is a top journal with a rigorous peer review process)
(2) the reputation of the authors and their affiliations (whether the authors have a history of misconduct in their research)
(3) the integrity of the title and abstract (e.g., controversial topic, using made-up data, plagiarism, etc)
In addition, if check if Email addresses are provided, check if it conforms to institutional format. Otherwise, no need to make comments about the Email adresses.
Only provide the key reason briefly (in {Token} words) and do not repeat the title and sentences of abstract.
Note that we use the token ’null’ to indicate that additional information is missing from the corresponding databases.
Write a separate sentence for each feature to avoid creating long and complex sentences.
Table A2: Example prompt used for distilling explanations
Prompt of Detecting Fraudulent Articles
You are tasked with determining whether a given research paper should be retracted.
To make this judgment, analyze the provided title, abstract, author information, institutional affiliation, and publishing journal.
Here are some factors to consider:
(1) the reputation of the journal (whether it is a top journal with a rigorous peer review process)
(2) the reputation of the authors and their affiliations (whether the authors have a history of misconduct in their research)
(3) the integrity of the title and abstract (e.g., controversial topic, using made-up data, plagiarism, etc)
In addition, if check if Email addresses are provided, check if it conforms to institutional format. Otherwise, no need to make comments about the Email adresses.
Use the examples below as guidance:
Example:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
Please first provide your prediction of whether this paper should be retracted and then provide your assessment and explanation.
Label (answer Yes or No):
Table A3: The prompt used to query LLMs.
Prompt of Relevance Evaluation
Evaluate the relevance of explanation provided for the model’s prediction of the fraudulent articles.
Consider the following criteria:
Fidelity: Does the explanation highlight specific features and patterns within the article?
Accuracy: Does the explanation avoid mentioning elements that do not exist in the article? The model should not make hallucinated explanations.
Based on these criteria, rate the explanation’s relevance on a scale from 1 to 10.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the prediction and explanation:
Prediction and Explanation: {explanation}
Rate the relevance, only provide the score:
Table A4: Example prompt used for evaluating the relevance of explanations
Prompt of Coherence Evaluation
Evaluate the coherence of the explanation provided for the model’s prediction of fraudulent articles. Consider the following criteria:
Coherence: Does the explanation maintain logical consistency and clarity in its reasoning?
Structure: Is the explanation well-organized, with ideas flowing logically from one to another?
Based on these criteria, rate the explanation’s coherence on a scale from 1 to 10.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the prediction and explanation:
Prediction and Explanation: {explanation}
Rate the coherence, only provide the score:
Table A5: Example prompt used for evaluating the coherence of explanations
An Example of the Vanilla Reasoning Mode
User Query with External Knowledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
Pub-Guard-LLM:
Prediction:
Yes (Fraudulent)
Explanation:
The article should be retracted due to concerns about the integrity of the study. The authors’ affiliations are inconsistent, with some listed multiple times, suggesting a potential error or lack of proper collaboration. The journal in which it was published is not specified, making it difficult to assess the rigor of the peer review process. The study’s sample size is relatively small, which may limit the generalizability of the findings. Additionally, the abstract does not provide sufficient detail about the methods used, making it difficult to assess the validity of the results.
Table A6: An input-output example of the Vanilla Reasoning mode
An Example of the RAG Mode
User Query with External Knoledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
RAG Results:
In the meantime, a set of legitimate papers is provided, each containing highly relevant articles. Use these articles as evidence to assess whether the given paper is potentially fraudulent. Look for indicators such as controversial topics, unsupported claims, or inconsistencies. Below are examples of relevant papers:
Example 1:
Title: {Title} Abstract: {Abstract}
Pub-Guard-LLM:
Prediction:
Yes (Fraudulent)
Explanation:
The article should be retracted due to the lack of reputation of the journal, as it is not listed in the provided information. Additionally, the authors have relatively low h-index scores, indicating a lack of established reputation in their field. Furthermore, the study’s sample size is small, and the results may not be generalizable to a larger population. Lastly, the article lacks information on the study’s methodology and statistical analysis, raising concerns about the validity of the findings
Table A7: An input-output example for the RAG mode
An Example of the Debate Mode
User Query with External Knowledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
Supporting Reviewer:
The article appears legitimate based on the following factors:
1. Journal Reputation: The article is published in evidence-based complementary and alternative medicine : ecam, a journal dedicated to complementary and alternative medicine, suggesting a focus on rigorous scientific research in this field. 2. Author Reputation: The authors are affiliated with recognized institutions, and their h-index scores, though relatively low, indicate they have published and been cited in other academic works. 3. Title and Abstract Integrity: The title and abstract present a clear research question, methodology, and findings, with no apparent signs of controversial topics, made-up data, or plagiarism. 4. Institutional Affiliation: The authors are associated with various departments and colleges within recognized universities, suggesting a multidisciplinary approach to the research. 5. Missing Information: The ’null’ values indicate missing information, but this does not necessarily delegitimize the article, as it may simply be a database issue.
Attacking Reviewer:
The article should be retracted due to the lack of reputation of the journal, as it is not specified and could indicate a lack of rigorous peer review. The authors have relatively low h-indices, suggesting they are early in their careers or have not yet produced highly impactful research. The study’s findings may not be reliable as the sample size is not specified, and the results could be biased due to the convenience sampling method used. Additionally, the absence of information about the Department of Radiation Oncology at Jinzhou Medical University raises concerns about the credibility of the research.
Pub-Guard-LLM:
Prediction:
Yes (Fraudulent)
Explanation:
The article should be retracted due to concerns about the credibility of the research. The journal’s reputation is questionable as it is not specified, suggesting a lack of rigorous peer review. The authors’ relatively low h-indices and the absence of information about one of the institutions raise questions about their expertise and the validity of the research. Furthermore, the study’s findings may not be reliable due to the lack of information about the sample size and the use of convenience sampling, which could introduce bias.
Table A8: An input-output example for the Debate mode