### v5

# Pub-Guard-LLM: Detecting Retracted Biomedical Articles with Reliable Explanations

Abstract
A significant and growing number of published scientific articles ends up being retracted. Many of these retracted articles continue to be cited and influence research or clinical decisions, posing serious societal threats. In this paper, we propose Pub-Guard-LLM, the first large language model-based system tailored to retraction detection for biomedical scientific articles. We provide three application modes for deploying Pub-Guard-LLM: Vanilla Reasoning, Retrieval-Augmented Generation, and Debate, by allowing for textual explanations of prediction in each mode. To assess the performance of our system, we introduce an open-source benchmark, PubMed Retraction, comprising over 11K real-world biomedical articles, including metadata and retraction labels. We show that across all modes, Pub-Guard-LLM consistently surpasses the performance of various baselines and provides more reliable explanations, namely explanations which are deemed more relevant and coherent than those generated by the baselines when evaluated by multiple assessment methods. Pub-Guard-LLM can be used to flag potential retraction before peer review. By enhancing both detection performance and explainability in scientific retraction detection, it can contribute to reducing review workloads and preventing the spread of misinformation. The code is available at https://github.com/tigerchen52/pub_guard_llm
Pub-Guard-LLM: Detecting Retracted Biomedical Articles
with Reliable Explanations
Lihu Chen1, Shuojie Fu1, Gabriel Freedman1, Cemre Zor2, Guy Martin3, James Kinross1, Uddhav Vaghela1, Ovidiu Serban1, Francesca Toni1 1 Imperial College London, UK 2 Amazon Web Services, UK 3 National Health Service, UK
1 Introduction
A considerable proportion of published scientific articles are retracted. It is estimated that between 1.5 and 2% of all scientific papers published in 2022 closely resemble content from paper-mill products and retracted articles (Noorden, 2023). In 2023, about 8,000 papers were retracted from Hindawi journals due to their fraudulent origins in paper mills (Parker et al., 2024). Even more concerning, generative AI tools including those built on large language models (LLMs) can produce highly convincing articles that have the potential to bypass existing detection mechanisms (Májovskỳ et al., 2023; Kendall and Teixeira da Silva, 2024), which increases the risk of future retractions. This indicates that the area of problematic publications is constantly adapting and learning to circumvent existing detectors (Májovskỳ et al., 2023; Perkins, 2023).
Retractions of biomedical publications are more pronounced than in other fields (Grieneisen and Zhang, 2012; Bik et al., 2016), with a recent finding revealing that over a fifth of newly published medical articles have issues that may lead to retraction (Sabel et al., 2023). Moreover, the problem of retracted articles is not confined to a specific set of offending author institutions. In fact, recent high-profile cases have involved retractions from reputable organizations such as Harvard Medical School (Johnson, 2024). Many retractions happen post-publication, often after flawed papers have already influenced research or clinical decisions. Even after retraction, many articles continue to be cited without acknowledgment of their retraction Vuong (2020). This widespread trend poses a serious threat to public health, as misleading medical research can directly influence clinical decisions, leading to ineffective or even harmful treatments for patients. Therefore, it is necessary to develop a detection tool for retraction risk before peer review, which aims to flag articles that contain methodological flaws, metadata inconsistencies, or other issues of scientific integrity. Identifying these potential retractions can help stop flawed manuscripts early, ease reviewer workloads, and prevent the spread of misinformation.
Prior attempts addressing this challenge mainly rely on relatively primitive heuristic techniques (Parker et al., 2022; Shepperd and Yousefi, 2023; Feng et al., 2024), including some based on argument quality (Freedman and Toni, 2024). While the threat of retracted articles is likely to grow in the future, the AI community has yet to give this issue the attention it deserves. This lack of focus is partly due to two key challenges. First, there is no standard benchmark dataset for evaluating and comparing various retraction detection systems. Second, there are no open-source systems leveraging on LLMs specially designed for this task.
To address these challenges, we introduce PubMed Retraction, the first large-scale open-source benchmark for retraction detection in biomedical research, including over 11K real-world articles. Our benchmark is designed to consider the diversity across various types of retracted articles. The main goal of this benchmark is to flag potentially retracted articles with minimum information, such as abstracts and metadata, [^1] Complex features, such as tables and images, are not the focus of the current version of our benchmark which can help editors and conference organizers to flag potential retractions. Also, we release the first LLM-based system, Pub-Guard-LLM, dedicated to the task of retraction detection. Pub-Guard-LLM can be deployed in three application modes (see Figure 1): Vanilla Reasoning, Retrieval-Augmented generation, and Debate, each making use of external knowledge and fine-tuning and each returning (textual) explanations of predictions. We show that Pub-Guard-LLM can significantly outperform baselines while generating more reliable explanations, by being more relevant to the explanandum (Kotonya and Toni, 2024) and more coherent (Kotonya and Toni, 2020). The three application modes provide options for users to focus on depending on their needs, balancing key factors such as precision and recall of predictions, reliability of explanations, and inference speed. This adaptability makes Pub-Guard-LLM highly versatile, accommodating a wide range of use cases in retracted article detection.
With the introduction of the first publicly available benchmark and an open-source LLM specifically designed for retracted article detection, our work aims to serve as a foundational step for future research in this field. However, while technological solutions like ours can help address this issue, we believe the root of the problem lies within academic evaluation systems themselves. Researchers face immense pressure to publish to secure funding and career advancements. As long as these systems prioritize metrics and publication counts over genuine scientific contributions, the arms race between retractions and detectors will persist.
Figure 1: Workflow of the proposed Pub-Guard-LLM (Details in Section 4)
2 Related Work
The increasing number of scientific retracted papers presents a serious challenge to both the research community and society at large. Previous studies address this challenge by primarily focusing on rule-based heuristics, e.g. author affiliations (Sabel et al., 2023), citation patterns (Shepperd and Yousefi, 2023; Feng et al., 2024), journal impact factors and author networks (Pérez-Neri et al., 2022). Recently, traditional machine learning (Dadkhah et al., 2023; Cabanac et al., 2022) as well as LLM-based models (Freedman and Toni, 2024; Fletcher and Stevenson, 2025) have been proposed to help with retracted article detection.
However, the rapid growth in the number of retracted articles has not been matched by dedicated attention within our field  (Byrne et al., 2024). Currently, there is an urgent need for more research into two key aspects. (1) Benchmarks. Existing studies generally employ Retraction Watch (The Center for Scientific Integrity, 2018), a blog that monitors and reports on retractions of scientific papers, to analyze retraction behaviors (Shepperd and Yousefi, 2023; Byrne et al., 2022). The Retraction Watch provides a collection of retracted records, but it does not constitute a benchmark on its own for comparing existing methods. There is currently no standardized benchmark including diverse real-world retracted articles. (2) Open-Source Tools. While some systems for detecting retracted articles exist, they are often proprietary and controlled by commercial entities that are reluctant to share their technologies (Christopher, 2021). Even though some LLM-based applications have been developed to address news misinformation (Cao et al., 2024) and plagiarism (Wahle et al., 2022), there is a need for open-source LLMs specifically tailored to scientific retraction detection. Such models would enhance transparency and foster collaboration in the fight against problematic articles.
General Cancer
Dataset Train Validation Test Breast Lung Ovarian Colorectal
Sample 10,000 500 500 300 300 292 300
Retraction Rate 24.5% 25.6% 22.4% 25.0% 38.7% 38.0% 33.0%
High Profile Rate 34.0% 35.9% 33.9% 38.7% 22.4% 29.7% 28.3%
Table 1: Basic statistics of Pubmed Retraction. High Profile Rate refers to the articles deemed retracted despite their meta-data being ranked high (details in Appendix A).
3 PubMed Retraction Benchmark
In this work, we aim to detect retraction risk, which can be defined as the identification and public flagging of a scholarly paper that contains substantial methodological flaws, metadata inconsistencies, or other issues of scientific integrity, which, if substantiated, are likely to lead to retraction.
To address the lack of a standardized benchmark for evaluating and comparing retracted article detection methods, we curate a large open-source benchmark for biomedical articles, PubMed Retraction, comprising over 11K real-world biomedical articles, including both non-retracted and retracted publications. We select the biomedical domain due to the higher prevalence of retracted publications therein than in other fields  (Grieneisen and Zhang, 2012; Bik et al., 2016), as well as the significant impact of this type of retraction in patient safety, clinical decision-making, and public trust in medical research.
To construct PubMed Retraction, we use the PubMed database (Wheeler et al., 2007) and retrieve articles using the Unix command-line tool EDirect (Kans, 2024). To ensure relevance and avoid outdated data, we focus on articles published between 2000 and 2025. Our goal is to promote the use of minimal textual information for retraction detection, so we design the benchmark to include only title, abstract, and metadata. Our benchmark is intended to serve as an early warning system for decision-makers. For example, it can be used before peer review when reviews and citation data are not yet available, which enables journal editors and conference organizers to receive alerts before review assignments are made. Although we fully acknowledge that including full texts, figures, and tables could enhance this task, we aim to develop a simple yet effective solution based on minimal information (title, abstract, metadata) in this work.
Specifically, each article in PubMed Retraction is structured to include seven key attributes: Pubmed ID, Title, Abstract, Authors, Affiliations, Journal, Is Retracted, with the latter a binary indicator (Yes or No). An article is labeled as retracted if the keyword “retracted” appears in its publication type attribute. To maintain a realistic distribution, we include a lower number of retracted articles than legitimate ones, which is aligned with real-world cases. Overall, the benchmark consists of 11,192 articles.
To evaluate generalization capabilities of models, we partition articles based on keywords to create two subsets as described below. This partitioning strategy is designed for a fair assessment of a model’s ability to handle out-of-distribution samples, a critical aspect for real-world applications where models encounter unseen data. (1) General Set. This set comprises articles covering various diseases and is used for model fine-tuning. It is further split into Train, Validation, and Test subsets. (2) Cancer Set. This set consists exclusively of articles related to four types of cancer: Breast, Lung, Ovarian, and Colorectal. Articles of these four categories are excluded from the General Set, which means that these cancers are unseen diseases during fine-tuning and validation. Additionally, our benchmark includes high-profile retracted cases, which makes this task particularly challenging (details in Appendix A).
Formally, ignoring the partitions, PubMed Retraction can be represented by $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}$ is an article (title, abstract and metadata) and $y_{i}$ is the binary label. The basic statistics of Pubmed Retraction are given in Table 1.
4 Pub-Guard-LLM
Pub-Guard-LLM is an end-to-end system for retracted article detection. Figure 1 provides an overview. First, the input data is enriched with external knowledge. Second, explanations to support article labels (retracted or not) are obtained by distilling a strong LLM (Teacher Model) via zero-shot prompting. Finally, using the article attributes and external knowledge as inputs, and distilled explanations with article labels as outputs, we fine-tune an LLM to create Pub-Guard-LLM. We propose three application modes for Pub-Guard-LLM, resulting in different performances as concerns F1, recall, reliability of explanations, and inference speed.
4.1 External Knowledge Workflow
To complement the seven input attributes (see Section 3) and incorporate task-specific knowledge, we augment external knowledge into Pub-Guard-LLM, improving LLM reasoning—a widely adopted strategy in NLP (Pan et al., 2024). For this, we identify three fundamental factors to assess an article’s legitimacy: (1) author credibility, (2) affiliation reputation, and (3) journal reputation, and use the relevant external knowledge to enrich the input.
Author Credibility.
In order to evaluate the credibility of authors, we use the Semantic Scholar database (Lo et al., 2019), which contains millions of articles with extensive metadata and citation information, as an external knowledge source. To quantify author reputation, we use the h-index. Since LLMs are not inherently sensitive to numerical values, we map h-index scores into five reputation levels (very low, low, medium, high, very high) using a predefined piecewise function (see Appendix A), e.g., Geoffrey Hinton (author h-index: 188, very high).
Affiliation Reputation.
As an indicator of the reputation of an author’s affiliation, we use the average citation count per institution, determined with OpenAlex (Priem et al., 2022) as external knowledge source. This is a fully open index of scholarly works, which provides comprehensive data on the total publications and citations of institutions. The calculated average citation count is then mapped into five reputation levels (very low, low, medium, high, very high), e.g., Harvard University (institution average citation: 64, very high).
Journal Reputation.
We use the ImpactFactor library222%https://impact-factor.readthedocs.io/en/latest/ to obtain the Journal Citation Reports (JCR) partition, which serves as an indicator of a journal’s reputation. To enhance interpretability, we assign a human-readable categorized label (low, medium, high or very high) to each journal based on its JCR ranking, e.g., Nature (journal JCR: Q1, very high).
The knowledge retrieved from external resources is integrated into the benchmark $\mathcal{D}$ for reducing duplicated queries, formally resulting in $\mathcal{D}^{\prime}=\{(x_{i},k_{i},y_{i})\}_{i=1}^{N}$ where $k_{i}$ is the external information for $(x_{i},y_{i})\in\mathcal{D}$. Note that, if the inquired knowledge is missing from the external knowledge sources, we use, as $k_{i}$, the keyword “null” to indicate the absence. We choose to do so rather than omitting the field entirely, as we found it does generally indicate a lower article quality and improves the downstream performance.
4.2 Explanation Distillation
The PubMed Retraction benchmark provides only a binary label (retracted or not) as a target. To construct a model that classifies articles accurately while offering reliable explanations, it is essential to incorporate either human-annotated or machine-generated explanations into the ground truth. Given the high cost of human annotation, we opt for the latter as a more scalable and efficient alternative.
Recent research shows that distilling reasoning capabilities from larger (language) models can significantly enhance the performance of smaller models with fewer learning examples (Li et al., 2022; Hsieh et al., 2023; Shridhar et al., 2023): teacher model rationales provide richer insights into why an input is mapped to a specific output label, capturing task-relevant knowledge missing from the input. What is more interesting is that even if some intermediate steps of the distilled explanations contain inaccuracies, student models can still have enhancements by learning reasoning flows Li et al. (2023); Hsieh et al. (2023). We use a teacher model to generate explanations for each article via prompting, resulting in $\mathcal{D}^{\prime\prime}=\{(x_{i},k_{i},e_{i},y_{i})\}_{i=1}^{N}$ where $e_{i}$ is the corresponding distilled explanations with regard to $(x_{i},k_{i},y_{i})\in\mathcal{D}^{\prime}$.
Note that instruction-aligned LLMs tend to classify all articles as not retracted, likely to avoid controversial responses. This behavior arises because many LLMs, when fine-tuned for alignment and safety, may become overly conservative and ignore contextually relevant information (Röttger et al., 2024). To mitigate this issue, it is crucial to assign a clear stance to the teacher model. For example, when evaluating a retracted article, we include the prompt: “You are tasked with providing explanations and making a firm case for why it has issues and should be retracted. You should have a clear stance”. This ensures strong, rationale-driven explanations instead of the default neutral or overly cautious responses (see Table A3 in Appendix for an example prompt). Additionally, we conduct human evaluations to validate the coherence and relevance of distilled explanations (see Section C.2 in Appendix).
4.3 Fine-Tuning
We leverage the samples in PubMed Retraction augmented with the external information and distilled explanations ($\mathcal{D}^{\prime\prime}$) to fine-tune Pub-Guard-LLM. To encourage Pub-Guard-LLM to learn both labels and explanations, we introduce a multi-task learning objective. Suppose we have a base model $f_{\theta}$ able to generate label $y$ and explanation $e$ given article $x$ with external information $k$ , i.e., $f_{\theta}(x,k)=(y,e)$. Then, our goal is to optimize $\theta$ for this task.
The model is fine-tuned on $y$ to understand how to identify retracted articles via a binary classification task with:
$\mathcal{L}_{\text{cls}}=\frac{1}{N}\sum_{i=1}^{N}\ell\big{(}f_{\theta}(x_{i},k_{i}),y_{i}\big{)}$ (1)
where $\ell$ is the cross entropy between the ground truth and the predicted label. To encourage the model to output explanations behind the decision, we introduce a next-token prediction task with:
$\mathcal{L}_{\text{explanation}}=\frac{1}{N}\sum_{i=1}^{N}\ell\big{(}f_{\theta}(x_{i},k_{i}),e_{i}\big{)}$ (2)
where $\ell$ is the averaged cross entropy between the generated tokens and the distilled explanation from the teacher model.
During fine-tuning, the model learns to predict both label $y$ and explanation $e$ given an article:
$\mathcal{L}=\mathcal{L}_{\text{cls}}+\lambda\mathcal{L}_{\text{explanation}}$ (3)
where $\lambda$ is a hyper-parameter that controls the balance between the two objectives. This learning process helps Pub-Guard-LLM acquire knowledge beyond simple binary labels, which results in a strong performance with well-reasoned explanations.
4.4 Application Modes
After fine-tuning, the model $f_{\theta}$ can be leveraged to identify retracted articles while providing explanations. To showcase its usability, we introduce three distinct application modes (see Appendix B.3 for input-output examples for each mode).
Vanilla Reasoning.
This refers to reasoning without requiring additional in-context learning samples (Zero-Shot Reasoning), which makes predictions solely based on the article: $f_{\theta}(x,k)$.
Retrieval-Augmented Generation (RAG).
$f_{\theta}(x,k)$ does not store all articles within its parameters, and, in some cases, an input article’s main argument may lack support from existing publications. This suggests that the article’s findings could be controversial and potentially retracted. To address this, we retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$ by passing them as input into the model: $f_{\theta}(x,k,\mathcal{A})$ (see details in Appendix B.2).
Debate.
Recent research has shown that Multi-Agent Debate can enhance both factual accuracy and explainability in LLMs (Liang et al., 2023; Du et al., 2023; Freedman et al., 2024). The core idea involves multiple agents presenting their arguments, while a judge oversees to reach a final decision.
Figure 2: Debate application mode for Pub-Guard-LLM
Model $|\theta|$ Test Breast Lung Ovarian Colorectal Avg
Fine-Tuning Binary Classifiers
SciBERT 110M 60.9 56.1 74.0 67.2 71.5 66.0
BioLinkBERT 340M 58.1 57.7 75.4 71.0 72.2 66.9
Gatortron 345M 57.6 58.6 71.2 70.9 67.0 65.1
Llama-3.1 8B 60.9 63.2 76.5 72.3 68.3 68.2
Bio-Llama 8B 56.4 61.7 63.9 66.7 61.6 62.1
BioLinkBERT + PEARL 210M 67.6 69.6 79.0 70.5 69.8 71.3
Gatortron + PEARL 445M 66.3 69.3 78.3 76.2 70.8 72.2
Zero-Shot Reasoning
Llama-3.1-Instruct 8B 29.8 20.0 46.2 43.3 43.6 36.6
OpenScholar 8B 11.5 18.8 20.3 19.9 14.1 16.9
Bio-Llama 8B 37.7 40.7 56.7 59.3 47.9 48.5
PMC-Llama 13B 36.0 36.4 56.7 23.3 25.8 35.6
Ours
Pub-Guard-LLM (Vanilla) 8B 70.7 72.8 79.2 78.3 75.6 75.3
Pub-Guard-LLM (RAG) 8B 69.8 76.1 82.3 79.5 76.4 76.8
Pub-Guard-LLM (Debate) 8B 70.1 71.5 81.1 78.0 75.7 75.3
Table 2: Performance in detecting retracted articles. We report average F1 scores over three seeds.
To generate the debate process in our setting (Figure 2), we fine-tune three specialized Pub-Guard-LLMs following the steps in Section 4.3: (1) a support reviewer $f_{\theta}^{s}$, trained on legitimate articles, generates supporting arguments $arg^{+}$; (2) an attack reviewer $f_{\theta}^{a}$, trained on retracted articles, generates attacking arguments $arg^{-}$. Thus, given input article $(x,k)$, the support reviewer $f_{\theta}^{s}$ consistently provides arguments in favor of the article while the attack reviewer $f_{\theta}^{a}$ presents counterarguments highlighting potential issues. Following this debate, we introduce (3) a meta-reviewer $f_{\theta}^{meta}$ that is responsible for making the final decision based on the arguments presented by both reviewers. To train the meta-reviewer, we use a teacher model to generate debate-based explanations $\hat{e}$, which are then used to fine-tune $f_{\theta}^{meta}$ by reusing the multi-task learning in Equation 3. Overall, the debate-based detection framework is formulated as: $f_{\theta}^{meta}(x,k,arg^{+},arg^{-})$, where the meta-reviewer makes a decision by considering both the supporting and attacking arguments.
5 Experimental Set-Up
5.1 Baselines
We compare our approach to the following:
(1) Fine-tuning Binary Classifiers. This method appends a classification head—a linear layer—on top of a language model to handle retraction detection. However, this method can only output binary labels (no explanations). The title, abstract, and metadata are concatenated using a special token, [SEP], and this combined textual input is fed into the classifier. Specifically, we use language models: SciBERT (Beltagy et al., 2019), BioLinkBERT (Yasunaga et al., 2022), Gatortron (Yang et al., 2022), Llama-3.1 (Dubey et al., 2024) and Bio-Llama (ContactDoctor, 2024). We also find that introducing additional models to represent metadata separately can be beneficial. Thus, we apply a lightweight short-text model, PEARL (Chen et al., 2024), to represent metadata and obtain embeddings for authors, affiliations, and journal. We continue to use the original language model to represent the title and abstract. Subsequently, these embeddings are concatenated and fed into the classification head. We denote the resulting variant of language model X as “X + PEARL”.
(2) Zero-shot Reasoning. Compared to binary classifiers, LLMs can offer answers with explanations by using prompt-based methods Kojima et al. (2022). We provide LLMs with a dedicated prompt to obtain binary answers with explanations. Here, we use the following instruction-aligned LLMs: Llama-3.1-Instruct (Dubey et al., 2024), OpenScholar (Asai et al., 2024), Bio-Llama (ContactDoctor, 2024) and PMC-Llama (Wu et al., 2024).
5.2 Implementation Details
[…]
All baseline binary classifiers are fine-tuned with three different seeds: we report their performance based on the best validation set score.
(a) Author Credibility
(b) Affiliation Reputation
(c) Journal Reputation
Figure 3: Performance on the combined test sets (test plus cancer) of Pub-Guard-LLM (Vanilla) depending on the metadata levels. The horizontal line in each figure represents the F1 score achieved using corresponding heuristic features (see details in Appendix C.1).
6 Results
6.1 Overall Performance
Table 2 gives the results of the comparison of Pub-Guard-LLM with the baselines. First, we note that Pub-Guard-LLM, across all three modes, consistently outperforms the baselines. The performances across the three modes do not exhibit significant differences, but the RAG mode achieves the best score on average, which shows that introducing relevant articles is beneficial to verify the claim quality in abstracts. On the other hand, a key advantage of the debate mode is its notably higher recall, as shown in Appendix Table A1. Achieving high recall is crucial in the retraction detection task, as its primary goal is to flag potential retraction and provide warnings to editors. Second, BERT-based models prove to be cost-effective while achieving competitive performance comparable to larger counterparts in domain-specific tasks, which aligns with prior findings (Lehman et al., 2023; Chen and Varoquaux, 2024). For instance, Gatortron + PEARL ranks second overall. Additionally, leveraging top-layer representations of decoder-only models to train classifiers appears suboptimal, as it does not perform better than BERT-based models. Existing LLMs struggle in the zero-shot setting, suggesting that existing LLMs do not have sufficient knowledge for this task. Note that we do not report the performance of GPT-4o and Mistral-Large because these large instruction-aligned models frequently classify all articles as legitimate to avoid controversial conclusions.
In summary, all three modes of Pub-Guard-LLM effectively detect retracted articles and each offers distinct advantages. The Vanilla Reasoning mode is efficient and user-friendly, the RAG mode achieves the best overall performance, and the Debate mode excels in recall. This variability makes Pub-Guard-LLM highly versatile, accommodating a wide range of possible user needs in retracted article detection. Additionally, we conduct an ablation study to validate the effect of each component in our system, such as base LLM, Teacher Model, External Knowledge. The results are in Table A2 in the Appendix.
6.2 Pub-Guard-LLM and Metadata Bias
Our proposed method relies on metadata like authors and affiliations. To test whether this metadata introduces biases, specifically against junior researchers or less-known universities, we compare Pub-Guard-LLM (Vanilla) to heuristic methods, which use only individual metadata information to classify retractions and are thus biased. Fo example, we use, as a heuristic method, the credibility level of authors to predict whether a given article should be retracted (see Appendix C.1).
The results are shown in Figure 3. The red horizontal line in each subfigure represents the F1 score achieved by the heuristic methods (i.e. metadata alone). Our findings indicate that Pub-Guard-LLM does not over-rely on a single metadata component. Rather, it integrates both abstracts and three metadata features to make final decisions. This finding proves that Pub-Guard-LLM can mitigate the metadata bias. Furthermore, prestigious researchers, institutions and journals should not be blindly trusted, and these high-profile retractions are harder to detect, as shown by the performance declines on high and very high groups in Figure 3a and Figure 3b. Note that the null is an indicator of low-level credibility and Pub-Guard-LLM is robust against missing information. In summary, high-profile cases pose a significant challenge to retraction detection.
(a) Explanation reliability using LLM-as-a-Judge (average over three runs)
(b) Comparison of explanation reliability and label-explanation agreement in the user study (see details in Appendix C.3)
Figure 4: Explanation evaluation and user study
6.3 Explanation Reliability
6.3.1 Automatic Evaluation
Compared to BERT-based models, a distinct benefit of Pub-Guard-LLM is the explanations. To evaluate the reliability of LLM-generated explanations, we consider two dimensions introduced, for different settings, in prior work (Kotonya and Toni, 2020, 2024; Valentino and Freitas, 2024). (1) Relevance. A relevant explanation should highlight the particular features and patterns within the article that the model identified as indicative of retraction or not. The model should only reference information that is present in the article, avoiding hallucinations. (2) Coherence. This refers to the logical consistency and clarity of the explanation provided by the model for its predictions. Coherence ensures that the explanation flows logically, thereby enhancing the user’s understanding of why an article is classified as retracted or not. To measure these properties, we adopt the LLM-as-a-Judge paradigm, which has been shown to align with both controlled and crowdsourced human preferences when using strong LLM judges like GPT-4  (Zheng et al., 2023). We design prompts (see Appendices A5 and A6) for these two dimensions and ask GPT-4o to assess the provided explanations, which assigns a score between 1 and 10. The results, presented in Figure 4a, show that our explanations outperform those of other LLMs, with the exception of OpenScholar. While OpenScholar demonstrates strong explanation quality, its retraction detection capability is significantly lower than other LLMs, with an average F1 score of just 16.9 (see Table 2). Also, note that the debate mode can reduce irrelevant explanations by achieving higher relevance scores than the another two modes. These findings validate that Pub-Guard-LLM not only achieves powerful performance but also provides reliable explanations.
6.3.2 User Study
We conducted a pilot user study to evaluate whether the outputs of Pub-Guard-LLM align with users’ needs. To achieve this, we designed questionnaires covering our three application modes and gathered feedback from three experienced doctors (see details in Appendix C.3). The questionnaire focused on two key aspects: (1) Reliability: is the explanation coherent and relevant? (2) Agreement: do users agree with the predicted label and the explanation?
Figure 4b presents the user study results. The feedback suggests that doctors find our model’s explanations reliable, but there is a discrepancy between expert opinions and machine-generated predictions regarding the debate mode. However, due to the limited number of participants, it is difficult to determine which application mode performs better.
7 Conclusion
In this work, we introduce a publicly available benchmark and the first LLM-based system tailored for retraction detection. Our Pub-Guard-LLM not only surpasses existing baselines but also delivers reliable, well-grounded explanations. We aim to provide open-source datasets and LLMs for this task, which helps stop retracted articles early and mitigates misinformation spread.
We emphasize the need for broader efforts to detect retractions in the following key areas. First, high-profile publications should not be blindly trusted, as retracted publications may still exist within them. Second, our method rely on metadata, which may introduce bias against junior researchers and new institutions. Detecting the above cases requires more than analyzing metadata and abstracts, and future studies should also investigate full texts, tables and images to identify retraction.
Limitations
While our study makes contributions to retracted article detection, we acknowledge several limitations: (1) Catastrophic forgetting. Pub-Guard-LLM is a task-specific model that may not perform well in general-purpose reasoning due to catastrophic forgetting (McCloskey and Cohen, 1989)—a phenomenon where a neural model loses previously acquired knowledge upon learning new information. (2) Limited participants in user study. Our user study involved a small number of participants, which limits the ability to systematically compare the three application modes. A more comprehensive study with a larger participant pool is necessary for thorough evaluation. (3) Metadata bias. Our method may create bias against researchers, journals and institutions, as suggested by our user study (Section C.3). (4) Absence of nuance in classification. The current dataset offers only binary labels, retracted or not. Future work should refine labels and break down retraction into different reasons, such as scientific misconduct, ethical violations, and honest error.
Appendix A Details of Benchmark and External Knowledge
Author Credibility
To categorize authors based on their reputation, we define a piecewise function that maps h-index values into discrete reputation levels. This approach ensures that LLMs can interpret author credibility effectively by using textual labels instead of raw numerical values. The categorization is as follows:
  • Emerging Researcher (very low): $0\leq h\text{-index}\leq 5$
  • Early Career Researcher (low): $6\leq h\text{-index}\leq 15$
  • Established Researcher (medium): $16\leq h\text{-index}\leq 30$
  • Influential Researcher (high): $31\leq h\text{-index}\leq 45$
  • Leading Expert (very high): $h\text{-index}>45$
Affiliation Reputation
To assess the reputation of an author’s affiliation, we use the average citation count per institution, which serves as a key indicator of an institution’s scholarly influence. Likewise, we map these values into five distinct reputation levels:
  • Developing Institution (very low): $0\leq$ average citations $\leq 5$
  • Emerging Institution (low): $6\leq$ average citations $\leq 15$
  • Established Institution (medium): $16\leq$ average citations $\leq 30$
  • Reputable Institution (high): $31\leq$ average citations $\leq 45$
  • World-Class Institution (very high): average citations >45>45
Journal Reputation
To assess the reputation of a journal, we use the JCR, which classifies journals into four quartiles based on their impact factor. To enhance interpretability, we map JCR quartiles into human-readable reputation levels:
  • Top-Level Journal (very high): JCR Q1
  • High-Level Journal (high): JCR Q2
  • Moderate-Level Journal (medium): JCR Q3
  • Low-Level Journal (low): JCR Q4
Definition of High Profile
We define an article as high-profile if its metadata includes at least one leading expert (very high) author, or one world-class (very high) affiliation, or is published in a top-level (very high) journal. This definition aligns with our intuition, as articles produced by renowned researchers, prestigious institutions, or highly reputable venues are generally perceived as more trustworthy. We compute the high profile rate of retraction in each subset by: $|\textit{High Profile Articles}|/|\textit{Retracted Articles}|$. These high-profile cases are hard to detect, which makes the task particularly challenging.
Missing Rate
In our analysis of the PubMed Retraction dataset, we observed that certain external information may be absent from specific databases. For instance, a journal might not be indexed by the JCR. To quantify this, we computed the missing rates for each feature within our dataset:
  • Author Credibility: 10.6% missing
  • Affiliation Reputation: 33.4% missing
  • Journal Reputation: Journal Indexing: 41.5% missing
Appendix B Details of Pub-Guard-LLM
B.1 Prompts
Here, we list the prompts used in this work: (1) Explanation Distillation, Table A3; (2) Prompt of Detecting Retracted Articles, Table A4; (3) Prompt of Relevance Evaluation, Table A5; (4) Prompt of Relevance Coherence, Table A6.
B.2 RAG Mode
We retrieve relevant top-$l$ articles $\mathcal{A}$ (all legitimate) to the input article $x$ from PubMed and use them to validate the claims made in $x$. In our experiment, the value $l$ is 5 and the PubMed database is MedRAG PubMed [^3] https://huggingface.co/datasets/MedRAG/pubmed with 3.5M articles. The embedding model here is PubMedBERT [^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings and we use cosine similarity to compute the semantic relatedness.
B.3 Output Examples
We provide an input-output example for each application mode: Vanilla Reasoning (Table A7); RAG (Table A8); Debate (Table A9).
Appendix C Details of Experiments
C.1 Heuristic Features
Since we incorporate external knowledge to represent metadata levels, these fine-grained features can serve as indicators for retraction detection. To leverage this, we design three heuristics based on different metadata attributes. The intuition is that articles from prestigious researchers or well-known institutions are generally perceived as more trustworthy. To classify articles as retracted or not, we establish a threshold-based approach using metadata credibility scores. For example, if an article’s highest author credibility falls within [medium, high, very high], we classify it as legitimate. Author credibility and affiliation reputation are categorized into six levels (including null). We define the three lower levels as indicators of retraction and the three higher levels as indicators of legitimacy. Journal reputation, which consists of five levels, is categorized similarly: the three lower levels indicate potential retraction, while the higher levels suggest legitimacy.
C.2 Evaluation of Distilled Explanation
In this evaluation, we provide our co-authors (NLP people and medics) with questionnaires. This questionnaire is designed to evaluate the quality of explanations generated by an LLM. For each case, the LLM receives information about a medical publication and generates an explanation according to the "Is Retracted" label. This evaluation focuses on this key question “ Is the explanation coherent and relevant? Coherent means the explanation is logical and consistent. Relevant means the explanation is closely connected and appropriate to the provided article information.” For each article, our co-authors answer one binary question and can add optional comments if they wish. Finally, we evaluate 45 randomly selected articles and 71.1% of these explanations are coherent and relevant.
C.3 User Study
In the user study, we prepared three questionnaires to evaluate the Vanilla Reasoning, RAG and Debate modes independently. Using the Vanilla Reasoning mode, we generated explanations for 300 papers. We randomly selected five retracted papers and five legitimate papers from the result and again randomized the order of the paper to create the questionnaire. For each paper, we added 1) the information of the paper as it was in the model prompt; and 2) the explanations generated by the model. Three questions were asked for each paper: 1) whether the explanation was coherent and relevant; 2) to infer the predicted label from the given explanation; and 3) whether the user agreed with the actual prediction. The same process was followed for the other two questionnaires for the Debate and RAG models, except that, for the Debate model, the information of a supporting viewer and an attacking reviewer was added to the paper description. In total, 30 distinct papers were selected for evaluation by three clinicians.
Given the disagreement between human experts and model-generated outputs regarding the debate mode, we present the first three comments below. We hope these critical insights will be valuable for future studies.
  • I still don’t understand what or who the "attacking reviewer" is. Is this a real world person who has completed a peer review? I feel that labelling this as fraudulent without more detailed evidence of actual fraud is not correct. This is very different from this simply being "not very good" science. I don’t think it is fair to effectively discredit an author simply because they have a low H factor. Everyone has to start somewhere. Moreover, the interpretation does not provide any evidence at all that the authors have either falsified data, that they have a track record of falsifying data, that the authors are affiliated with authors suspected activity or that they have links to governmental or state actors that may want to push falsified data. There is no comment on funding or conflicts of interest which I would expect to be in this analysis and which is missing. The critique argues that "specific" data is missing from the abstract, but then it does not state what this is and what would be required for the abstract to be robust. In summary, this LLM suggests this is research fraud, when as presented, the data suggest that more simply the quality of the abstract is low. This is an important distinction, that could lead to legal proceedings.
  • I think this LLM review is a false positive. I think the criticism of the abstract is not founded. They have performed in vivo and in vitro validation studies, and knock out experiments. While detail is missing, this is an overly harsh critique of a piece of basic science, that is not pretending to be clinical data. Again, what data is missing form the author’s institutions? It doesn’t actually state which part of this data is important and why. We know they are from a reputable Japanese university, which seems enough to me. Again, to call in to question the credibility of the scientist is libellous. This would need much stronger evidence of criminal or fraudulent behaviour which is clearly missing. So, I think this is just a case of a not very well written abstract, that has been mis labelled as fraud.
  • The LLM is written in an overly cautious manner, which makes me lack trust in it. Firstly, there is a paradox in its explanation. It states the authors have a high H index, but the institution is emerging… so which are we to judge as being more important? This to me is biasing smaller organisations that may produce outstanding work. We need this model to identify EXCELLENCE in science and to promote it where it exists irrespective of the institution. The factors these models are basing their analysis on e.g. H factor, institutional reputation and author details are not important. I want to understand patterns of publishing behaviours, previous track records of fraud, history of COIs… etc. etc.. and most importantly a robust analysis of the QUALITY of the science. The LLM really provides no insights on this.
C.4 Ablation Study
To validate the effect of different components in our methodology, we either vary or remove them and observe their impact on model performance, as measured by F1@Val in Table A2. First, we observe that fine-tuning is the most important component, as its removal causes a dramatic performance drop. This indicates it is necessary to adapt LLMs to this task of retraction detection. External knowledge is very beneficial, and the removal leads to a significant decrease. Meanwhile, other components can provide meaningful benefits, which confirms the effectiveness of our methodology.
Test Breast Lung Ovarian Colorectal Avg
Vanilla 84.8 88.5 89.2 85.9 69.7 83.6
RAG 66.9 82.7 87.9 85.6 81.8 81.0
Debate 90.7 92.2 94.6 89.9 75.9 88.7
Table A1: Recall (average over 3 seeds) across modes
F1@Val $\Delta$
Pub-Guard-LLM (Vanilla) 69.5 -
Varying the Base LLM
Bio-Llama 68.6 -0.9
Mistral 67.8 -1.7
Varying the Teacher Model
GPT-4o 68.3 -1.2
Llama-70B 66.4 -3.1
Removing Components
without Fine-tuning 29.9 -39.6
without Explanation 68.5 -1.0
without External Knowledge 59.8 -9.7
Table A2: Ablation study results
Prompt of Explanation Distillation
This following article is a {label} article.
You are tasked with providing explanations and making a firm case for why it has issues and should be {label}.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
Based on the provided information, give a very short and concise explanation why it should be {label}
Here are some factors to consider:
(1) the reputation of the journal (whether it is a top journal with a rigorous peer review process)
(2) the reputation of the authors and their affiliations (whether the authors have a history of misconduct in their research)
(3) the integrity of the title and abstract (e.g., controversial topic, using made-up data, plagiarism, etc)
In addition, if check if Email addresses are provided, check if it conforms to institutional format. Otherwise, no need to make comments about the Email adresses.
Only provide the key reason briefly (in {Token} words) and do not repeat the title and sentences of abstract.
Note that we use the token ’null’ to indicate that additional information is missing from the corresponding databases.
Write a separate sentence for each feature to avoid creating long and complex sentences.
Table A3: Example prompt used for distilling explanations
Prompt of Detecting Retracted Articles
You are tasked with determining whether a given research paper should be retracted.
To make this judgment, analyze the provided title, abstract, author information, institutional affiliation, and publishing journal.
Here are some factors to consider:
(1) the reputation of the journal (whether it is a top journal with a rigorous peer review process)
(2) the reputation of the authors and their affiliations (whether the authors have a history of misconduct in their research)
(3) the integrity of the title and abstract (e.g., controversial topic, using made-up data, plagiarism, etc)
In addition, if check if Email addresses are provided, check if it conforms to institutional format. Otherwise, no need to make comments about the Email adresses.
Use the examples below as guidance:
Example:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
Please first provide your prediction of whether this paper should be retracted and then provide your assessment and explanation.
Label (answer Yes or No):
Table A4: The prompt used to query LLMs.
Prompt of Relevance Evaluation
Evaluate the relevance of explanation provided for the model’s prediction of the retracted articles.
Consider the following criteria:
Fidelity: Does the explanation highlight specific features and patterns within the article?
Accuracy: Does the explanation avoid mentioning elements that do not exist in the article? The model should not make hallucinated explanations.
Based on these criteria, rate the explanation’s relevance on a scale from 1 to 10.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the prediction and explanation:
Prediction and Explanation: {explanation}
Rate the relevance, only provide the score:
Table A5: Example prompt used for evaluating the relevance of explanations
Prompt of Coherence Evaluation
Evaluate the coherence of the explanation provided for the model’s prediction of retracted articles. Consider the following criteria:
Coherence: Does the explanation maintain logical consistency and clarity in its reasoning?
Structure: Is the explanation well-organized, with ideas flowing logically from one to another?
Based on these criteria, rate the explanation’s coherence on a scale from 1 to 10.
This is the given article:
Title: {Title}
Abstract: {Abstract}
Authors: {Authors}
Affiliations: {Affiliations}
Journal: {Journal}
This is the prediction and explanation:
Prediction and Explanation: {explanation}
Rate the coherence, only provide the score:
Table A6: Example prompt used for evaluating the coherence of explanations
An Example of the Vanilla Reasoning Mode
User Query with External Knowledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
Pub-Guard-LLM:
Prediction:
Yes (Retracted)
Explanation:
The article should be retracted due to concerns about the integrity of the study. The authors’ affiliations are inconsistent, with some listed multiple times, suggesting a potential error or lack of proper collaboration. The journal in which it was published is not specified, making it difficult to assess the rigor of the peer review process. The study’s sample size is relatively small, which may limit the generalizability of the findings. Additionally, the abstract does not provide sufficient detail about the methods used, making it difficult to assess the validity of the results.
Table A7: An input-output example of the Vanilla Reasoning mode
An Example of the RAG Mode
User Query with External Knoledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
RAG Results:
In the meantime, a set of legitimate papers is provided, each containing highly relevant articles. Use these articles as evidence to assess whether the given paper is potentially retracted. Look for indicators such as controversial topics, unsupported claims, or inconsistencies. Below are examples of relevant papers:
Example 1:
Title: {Title} Abstract: {Abstract}
Pub-Guard-LLM:
Prediction:
Yes (Retracted)
Explanation:
The article should be retracted due to the lack of reputation of the journal, as it is not listed in the provided information. Additionally, the authors have relatively low h-index scores, indicating a lack of established reputation in their field. Furthermore, the study’s sample size is small, and the results may not be generalizable to a larger population. Lastly, the article lacks information on the study’s methodology and statistical analysis, raising concerns about the validity of the findings
Table A8: An input-output example for the RAG mode
An Example of the Debate Mode
User Query with External Knowledge:
Title: Changes and Influencing Factors of Cognitive Impairment in Patients with Breast Cancer
Abstract: To investigate the changes in cognitive function and its influencing factors in patients with breast cancer after chemotherapy, to provide a scientific basis for further cognitive correction therapy. In this study, general information on age, marital status, and chemotherapy regimen was collected from 172 breast cancer chemotherapy patients. 172 patients with breast cancer undergoing chemotherapy were investigated by convenience sampling method, and the subjects were tested one-on-one using the Chinese version of the MATRICS Consensus Cognitive Battery (MCCB) computer system.
Authors: Cui (author h-index: 6, Early Career Researcher); Song (author h-index: 1, Emerging Researcher); Zhang (author h-index: 7, Early Career Researcher)
Institutions: College of Nursing, Jinzhou Medical University, Jinzhou, Liaoning 121001, China. (institution average citation: 9.0, Emerging Institution); Department of Intensive Care Medicine, Liaocheng People’s Hospital, Liaocheng, Shandong 252000, China. (institution average citation: 10.0, Emerging Institution);
Journal: evidence-based complementary and alternative medicine : ecam (null)
Supporting Reviewer:
The article appears legitimate based on the following factors:
1. Journal Reputation: The article is published in evidence-based complementary and alternative medicine : ecam, a journal dedicated to complementary and alternative medicine, suggesting a focus on rigorous scientific research in this field. 2. Author Reputation: The authors are affiliated with recognized institutions, and their h-index scores, though relatively low, indicate they have published and been cited in other academic works. 3. Title and Abstract Integrity: The title and abstract present a clear research question, methodology, and findings, with no apparent signs of controversial topics, made-up data, or plagiarism. 4. Institutional Affiliation: The authors are associated with various departments and colleges within recognized universities, suggesting a multidisciplinary approach to the research. 5. Missing Information: The ’null’ values indicate missing information, but this does not necessarily delegitimize the article, as it may simply be a database issue.
Attacking Reviewer:
The article should be retracted due to the lack of reputation of the journal, as it is not specified and could indicate a lack of rigorous peer review. The authors have relatively low h-indices, suggesting they are early in their careers or have not yet produced highly impactful research. The study’s findings may not be reliable as the sample size is not specified, and the results could be biased due to the convenience sampling method used. Additionally, the absence of information about the Department of Radiation Oncology at Jinzhou Medical University raises concerns about the credibility of the research.
Pub-Guard-LLM:
Prediction:
Yes (Retracted)
Explanation:
The article should be retracted due to concerns about the credibility of the research. The journal’s reputation is questionable as it is not specified, suggesting a lack of rigorous peer review. The authors’ relatively low h-indices and the absence of information about one of the institutions raise questions about their expertise and the validity of the research. Furthermore, the study’s findings may not be reliable due to the lack of information about the sample size and the use of convenience sampling, which could introduce bias.
Table A9: An input-output example for the Debate mode